# Two-Electron Correlations in the Metallic Electron Gas

### v1

@@ Abstract
Furthermore, we calculate the full e-e scattering amplitude with high precision across the entire Fermi surface. Crucially, these results serve as the foundation for our proposal of a robust effective interaction: a charge-based KO model within the local density approximation (LDA), supplemented by an s-wave correction (sKO+). This ansatz effectively resolves the discrepancies found in widely used approximations, such as the RPA and the KO interaction. Notably, our sKO+ ansatz offers a consistent description of electron correlations across all energy scales, extending the high precision rigorously established on the Fermi surface.
@@
Figure 3 displays the angle-resolved Landau quasiparticle interaction $F_{\sigma\sigma^{\prime}}(\theta)$ versus $\cos\theta$ at $r_{s}=1$ and $5$ for both parallel ($\uparrow\uparrow$) and anti-parallel spin ($\uparrow\downarrow$) configurations. The bare exchange Coulomb interaction $W_{\rm Ex}(\mathbf{k}_{1},\mathbf{k}_{2})=-N_{F}v(|\mathbf{k}_{1}-\mathbf{k}_{2}|)$ is shown for comparison. In the parallel spin channel, many-body screening substantially modifies the quasiparticle interaction from the bare exchange. At $r_{s}=1$, $F_{\uparrow\uparrow}(\theta)$ deviates from the bare exchange primarily at small angles with rapid variation as $\theta\to\pi$, while at $r_{s}=5$, deviations occur across all angles with a smoother angular dependence. This evolution indicates increasing locality of the quasiparticle interaction and enhanced screening at larger $r_{s}$ values. For the anti-parallel spin channel, where the bare two-electron short-range interaction vanishes, the quasiparticle interaction emerges entirely from many-body effects. Notably, $F_{\uparrow\downarrow}(\theta)$ exhibits minima at $\theta=\pi$ for both $r_{s}$ values, where the incoming electron’s momentum and spin are opposite, potentially relating to Cooper pairing instability from dynamic screening potentieffects [55].
@@
Figure 4 presents the dimensionless two-electron scattering amplitude, $A^{\sigma\sigma^{\prime}}(\theta,\phi)=N_{F}^{*}a_{\sigma\sigma^{\prime}}(\theta,\phi)$, for the uniform electron gas. We analyse the forward-scattering ($\phi=0$) and Cooper ($\theta=\pi$) channels at $r_{s}=1$ and 5 for $T/T_{F}=0.025$, and include $T/T_{F}=0.05$ for $r_{s}=5$. Comparison of these datasets demonstrates that our results have reached the zero-temperature limit within error bars, except for the Cooper channel divergence caused by the Tolmachev sigularity [55]. Crucially, as this singularity is confined to a phase space of measure zero, it has a negligible impact on Fermi-surface integrals in general. Based on our numerical results, we conjecture an effective interaction (sKO+), which is more accurate than the common effective interaction, for example, the RPA and KO. This ansatz is based on the charge KO interaction within the LDA, $f^{+}_{\rm XC}(q)\approx f_{\rm XC}^{+}(0)$($f_{\rm XC}^{+}$ below), supplemented by an s-wave correction $\delta R$ for the antiparallel spin configuration:
@@
Disccusion.—In this work, we present high-precision ab initio calculations of the four-point vertex function for the 3D uniform electron gas using our developed VDMC framework. We extract Landau parameters that validate the local density approximation and reveal how the dominant $F_{0}^{s}$ parameter drives the system from underscreening toward overscreening as electron density decreases. Our analysis of electron-electron scattering amplitudes on the Fermi surface enables us to establish a simple yet accurate effective model, the sKO+ ansatz, which is based on the charge KO interaction in LDA with a small s-wave correction. The sufficiency of such short-range corrections parallels recent findings that imperfect screening at short distances necessitates such corrections to capture the physics of the low-density regime. [61]. The excellent agreement between experimental thermal resistivity measurements, our numerical data, and this effective model demonstrates that our results provide a robust foundation for first-principles transport calculations for the electron liquids in simple metals. Looking forward, our methods offer crucial numerical benchmarks for emerging experimental techniques that probe two-particle correlations in materials [15, 16, 17]. From the theoretical point of view, the sKO+ effective interaction will inspire us to construct a better low-energy effective field theory for the fermionic system, consistent with recent insights into the critical role of short-range interactions in low-density metals.
Acknowledgement—K. C. was supported by the National Key Research and Development Program of China, Grant No. 2024YFA1408604, and the National Natural Science Foundation of China under Grants No. 12047503. Z.L, P.H., and Y.D. were supported by the National Natural Science Foundation of China (under Grant No. 12275263), the Innovation Program for Quantum Science and Technology (under grant No. 2021ZD0301900), the Natural Science Foundation of Fujian Province of China (under Grant No. 2023J02032).

### v2

@@ Abstract
Furthermore, we calculate the full e-e scattering amplitude with high precision across the entire Fermi surface. Crucially, these results serve as the foundation for our proposal of a robust effective interaction: a charge-based KO model within the local density approximation (LDA), supplemented by an s-wave correction for the antiparallel-spin channel (sKO+). The resulting effective interaction is given by
This ansatz effectively resolves the discrepancies found in widely used approximations, such as the RPA and the KO interaction. Notably, our sKO+ ansatz offers a consistent description of electron correlations across all energy scales, extending the high precision rigorously established on the Fermi surface.
@@
Figure 3 displays the angle-resolved Landau quasiparticle interaction $F_{\sigma\sigma^{\prime}}(\theta)$ versus $\cos\theta$ at $r_{s}=1$ and $5$ for both parallel ($\uparrow\uparrow$) and anti-parallel spin ($\uparrow\downarrow$) configurations. The bare exchange Coulomb interaction $W_{\rm Ex}(\mathbf{k}_{1},\mathbf{k}_{2})=-N_{F}v(|\mathbf{k}_{1}-\mathbf{k}_{2}|)$ is shown for comparison. In the parallel spin channel, many-body screening substantially modifies the quasiparticle interaction from the bare exchange. At $r_{s}=1$, $F_{\uparrow\uparrow}(\theta)$ deviates from the bare exchange primarily at small angles with rapid variation as $\theta\to 0$, while at $r_{s}=5$, deviations occur across all angles with a smoother angular dependence. This evolution indicates increasing locality of the quasiparticle interaction and enhanced screening at larger $r_{s}$ values. For the anti-parallel spin channel, where the bare two-electron short-range interaction vanishes, the quasiparticle interaction emerges entirely from many-body effects. Notably, $F_{\uparrow\downarrow}(\theta)$ exhibits minima at $\theta=\pi$ for both $r_{s}$ values, where the incoming electron’s momentum and spin are opposite, potentially relating to Cooper pairing instability from dynamic screening potentieffects [6].
@@
Figure 4 presents the dimensionless two-electron scattering amplitude, $A^{\sigma\sigma^{\prime}}(\theta,\phi)=N_{F}^{*}a_{\sigma\sigma^{\prime}}(\theta,\phi)$, for the uniform electron gas. We analyse the forward-scattering ($\phi=0$) and Cooper ($\theta=\pi$) channels at $r_{s}=1$ and 5 for $T/T_{F}=0.025$, and include $T/T_{F}=0.05$ for $r_{s}=5$. Comparison of these datasets demonstrates that our results have reached the zero-temperature limit within error bars, except for the Cooper channel divergence caused by the Tolmachev sigularity [6]. Crucially, as this singularity is confined to a phase space of measure zero, it has a negligible impact on Fermi-surface integrals in general. Based on our numerical results, we conjecture an effective interaction (sKO+), which is more accurate than the common effective interaction, for example, the RPA and KO. This ansatz is based on the charge KO interaction within the LDA, $f^{+}_{\rm XC}(q)\approx f_{\rm XC}^{+}(0)$($f_{\rm XC}^{+}$ below), supplemented by an s-wave correction $\delta R$ for the antiparallel spin configuration as presented in Eqs. (1),(2).
@@
Disccusion.—In this work, we present high-precision ab initio calculations of the four-point vertex function for the 3D uniform electron gas using our developed VDMC framework. We extract Landau parameters that validate the local density approximation and reveal how the dominant $F_{0}^{s}$ parameter drives the system from underscreening toward overscreening as electron density decreases. Our analysis of electron-electron scattering amplitudes on the Fermi surface enables us to establish a simple yet accurate effective model, the sKO+ ansatz, which is based on the charge KO interaction in LDA with a small s-wave correction. The sufficiency of such short-range corrections parallels recent findings that imperfect screening at short distances necessitates such corrections to capture the physics of the low-density regime. [27]. The excellent agreement between experimental thermal resistivity measurements, our numerical data, and this effective model demonstrates that our results provide a robust foundation for first-principles transport calculations for the electron liquids in simple metals. Looking forward, our methods offer crucial numerical benchmarks for emerging experimental techniques that probe two-particle correlations in materials [13, 51, 15]. Moreover, our work paves the way for next-generation ab initio methods targeting strongly correlated materials, where downfolding the UEG vertex function is essential for constructing accurate effective Hamiltonians. From the theoretical point of view, the sKO+ effective interaction will inspire us to construct a better low-energy effective field theory for the fermionic system, consistent with recent insights into the critical role of short-range interactions in low-density metals.
Acknowledgement—K.C. extends sincere thanks to Boris Svistunov and Carl A. Kukkonen for their valuable discussions. K. C. was supported by the National Key Research and Development Program of China, Grant No. 2024YFA1408604, the National Natural Science Foundation of China under Grants No. 12474245 and No. 12447103, and the GHfund A(202407010637). Z.L, P.H., and Y.D. were supported by the National Natural Science Foundation of China (under Grant No. 12275263), the Innovation Program for Quantum Science and Technology (under grant No. 2021ZD0301900), the Natural Science Foundation of Fujian Province of China (under Grant No. 2023J02032).